# Are Large Language Models Good Statisticians?

### v1

@@ D.1 Examples of Error Taxonomy
D.2 Error Analysis for Introducing CoT
@@
As presented in Table 10, the 0-shot overall performance of smaller LLMs, namely LLaMA-2-7b, LLaMA-3-8b, and LLaMA-3-8b-Instruct, experiences a slight decline following the introduction of CoT. This performance degradation is most notable in descriptive statistics tasks. However, the negative impact of CoT on 1-shot learning experiments is mitigated, resulting in nearly identical overall performance between one-shot and one-shot-CoT experiments. In contrast, for larger models such as LLaMA-2-13b and GPT models, the influence of CoT is negligible.
@@
As illustrated in Figure 9, the introduction of CoT in smaller models including LLaMA-2-7b, LLaMA-3-8b, and LLaMA-3-8b-Instruct, tends to slightly increase the incidence of invalid answers and column selection errors, leading to a marginal decline in overall performance. However, the impact of CoT on error distribution in 1-shot experiments or on LLaMA-2-13b and GPT models is not prominent.
@@
D.3 Details of Statistical Task Confusion Error
@@
Appendix E Qualitative Analysis
@@ E.2 Failure Examples
E.3 Other Notable Cases
@@ Appendix F Limitations
Appendix G Ethic Statement
This paper evaluates the capabilities of LLMs in our task and reveals the different strengths of LLMs and humans. This is not intended to provoke anxiety, but rather to gain a better understanding of LLMs’ abilities and to promote the development of LLMs. Additionally, it aims to foster discussions on how humans and LLMs can complement each other, and how humans can better utilize LLM tools in this era.
@@ Appendix G Ethic Statement
Appendix H Datasheet for StatQA
@@ H.1 Motivation
The creation of StatQA is funded by the Red Bird Grant and Research Grant from The Hong Kong University of Science and Technology (Guangzhou).
@@ H.3 Collection Process
H.4 Preprocessing, Cleaning, and/or Labeling

### v2

@@ E.1 Examples of Error Taxonomy
E.2 Error Analysis for Introducing CoT
@@
As presented in Table 12, the 0-shot overall performance of smaller LLMs, namely LLaMA-2-7b, LLaMA-3-8b, and LLaMA-3-8b-Instruct, experiences a slight decline following the introduction of CoT. This performance degradation is most notable in descriptive statistics tasks. However, the negative impact of CoT on 1-shot learning experiments is mitigated, resulting in nearly identical overall performance between one-shot and one-shot-CoT experiments. In contrast, for larger models such as LLaMA-2-13b and GPT models, the influence of CoT is negligible. As illustrated in Figure 9, the introduction of CoT in smaller models including LLaMA-2-7b, LLaMA-3-8b, and LLaMA-3-8b-Instruct, tends to slightly increase the incidence of invalid answers and column selection errors, leading to a marginal decline in overall performance. However, the impact of CoT on error distribution in 1-shot experiments or on LLaMA-2-13b and GPT models is not prominent.
@@
E.3 Error Analysis for Introducing Domain Knowledge
As reported and analyzed in Section 3.2, the introduction of domain knowledge (DK) significantly boosts the overall performance of LLMs, particularly with larger proprietary LLMs like GPT models. However, as also shown in Table 2, it adversely affects some models’ accuracy in correlation analysis (CA) and descriptive statistics (DS) tasks.
To further investigate, we also conduct an error analysis comparing 1-shot and 1-shot+DK approaches in the CA task, as depicted in Figure 10. It reveals a reduction in applicability errors across most LLMs, yet an increase in column selection errors and statistical task confusion among certain LLMs (i.e., LLaMA-2-13b, LLaMA-3-8b-Instruct, GPT-4, GPT-4o) following the introduction of DK. Conversely, DK tends to positively influence the performance of more complex and less common tasks like DCT, CTT, and VT, thus contributing positively to overall performance.
We propose that the observed phenomenon may be attributed to the following potential factors: (1) Since DCT, CTT, and VT tasks are less commonly seen compared to CA and DS, LLMs are less involved in their pre-training data, making them relatively unfamiliar with these tasks, so the introduction of DK obtains substantial improvement; (2) For relatively simple and straightforward tasks (CA and DS), where LLMs possess some familiarity, the introduction of DK possibly leads to information obfuscation, resulting in varying levels of performance degradation on simpler subtasks. Deeper reasons may stem from various processes of LLMs’ pre-training and the underlying mechanics of inference, which represents a valuable avenue for future research.
Figure 10: Distribution of Error Categories Across with/without DK on Correlation Analysis Task.
E.4 Details of Statistical Task Confusion Error
@@
Appendix F Qualitative Analysis
@@ F.2 Failure Examples
F.3 Other Notable Cases
@@ Appendix G Limitations
Appendix H Ethic Statement
This paper evaluates the capabilities of LLMs in our task and reveals the different strengths of LLMs and humans. This is not intended to provoke anxiety, but rather to gain a better understanding of LLMs’ abilities and to promote the development of LLMs. This study aims to foster discussions on how humans and LLMs can complement each other, and how humans can better utilize LLM tools in this era. Additionally, given the risks associated with LLMs producing erroneous or even toxic information, we advise readers to approach content generated by LLMs with caution.
@@ Appendix H Ethic Statement
Appendix I Datasheet for StatQA
@@ I.1 Motivation
The creation of StatQA is supported by NSF of China (62402409), Guangdong Basic and Applied Basic Research Foundation (2023A1515110545), CCF-Huawei Populus Grove Fund (CCF-HuaweiDB202403), and the Red Bird Grant, Research Grant from The Hong Kong University of Science and Technology (Guangzhou).
@@ I.3 Collection Process
I.4 Preprocessing, Cleaning, and/or Labeling